C = 50 + 0.80 (Y – T)
I = 200
G = 100
Find out the equilibrium level of income.
Suppose G increases to 125 what is the new equilibrium level of income?
What level of G is needed to achieve a target income of 200?
1
Expert's answer
2020-05-25T09:45:35-0400

(a) we know for the equilibrium condition,

Total income = Total expenditure

Y= C+I+G

here symbols have their usual meaning

Y= 50 + 0.8(Y-T) +I+G

Y= 50 + 0.8Y - 0.8 T+200+100

0.2 Y+0.8T=350

Y+0.4T= 1750, here T is income tax and Y is the income


(b) when G increases to 125 then,

New equilibrium level will be

Y= 50 + 0.8Y - 0.8 T+200+125

0.2 Y+0.8T=375

Y+0.4T = 1875


(c) For target income to be achieve as 200

Y= 50 + 0.8Y - 0.8 T+200+G

0.2 Y +0.8T = 250 + G

40 +0.8 T = 250 +G

G= -210 + 0.8 T